sarah was born when her father was 24 years old.After how long will her father be three times her age? shedunankai@gmail.com
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52788)

.
Let x = "after how long" years.


Then your equation is

3x = x + 24


2x = 24  ====>  x = 24/2 = 12.


In 12 years, Sarah is 12, while her father is 12 + 24 = 36, and 36 is 3 times 12.

Solved.


Answer by MathTherapy(10552)

sarah was born when her father was 24 years old.After how long will her father be three times her age? shedunankai@gmail.com
Depends on his or her current age.